ter 2 Assignment Use the bar graph from the World View. Comparative Output (GDP) to answer three questions. us In 2015, the worlds total output (real GDP) was roughly $80 trillion. What percentage of this total was produced Instructions: Enter your responses rounded to two decimal places a. By the United States and China? b. By Haiti and Ethiopia? Prev 1 of 12 Next>
Answer #1

Percentage of total produced in US and China = (US GDP + China's GDP)/Total GDP x 100

Percentage of total produced in US and China = (18 + 11)/80 x 100 = 36.25%

Similarly,

Percentage of total produced in Haiti and Ethopia = (0.01 + 0.06)/80 x 100 = 0.09%